40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Nov 14 16:26:31 2010, jenne, valera, Update, Electronics, SRM side OSEM noise with no magnet SRM-SD-Current-NoMagnet.pdf
    Reply  Mon Nov 15 04:57:10 2010, rana, Update, Electronics, SRM side OSEM noise with no magnet 
Message ID: 3916     Entry time: Sun Nov 14 16:26:31 2010     Reply to this: 3918
Author: jenne, valera 
Type: Update 
Category: Electronics 
Subject: SRM side OSEM noise with no magnet 

We realized that the SRM sensors are connected to the readout but just sitting on the BS in vacuum table with no magnets and therefore no shadows in them. We swapped the inputs to the SRM and PRM satellite boxes to use the higher transimpedance gain of the PRM side sensor. The attached plot shows the current spectrum in this configuration. The PD readback voltage was 9.5 V. Since this is close to the rail we put a slightly higher voltage into the AA of this channel to test that we can read out more ADC counts to make sure we are not saturating. The margin was 15800 vs 15400 counts with p-p of 5 counts on the dataviewer 1 second trend. We returned all cables to nominal configuration.

The calibration from A to m is 59 uA/1 mm.

Attachment 1: SRM-SD-Current-NoMagnet.pdf  171 kB  Uploaded Sun Nov 14 17:56:38 2010  | Hide | Hide all
SRM-SD-Current-NoMagnet.pdf
ELOG V3.1.3-